Sol Summary – February 13th

Crew 292 Sol Summary Report 13-02-2024

Sol: Sol # 9

Summary Title: Another remarkable day for Mangalyatri Crew 292

Author’s name: Annalea Beattie

Mission Status: Nominal

Sol Summary Activity: This morning all crew ate blueberry pancakes together and over breakfast we talked again about the future Indian Mars Desert Research Station. This time we discussed how we conserve resources here and then, our waste here, what happens to it and how other remote facilities everywhere manage their refuse. We decided to create a combined research project for Ladakh that focuses on waste disposal and see if we can create some interest and support from engineers and scientists and students in India. And each of our crew will contribute from their own position. For instance, Rajvi Patel, our engineer, works on Boeing aircrafts so she can investigate best practice for how they dispose of lithium in their batteries.

After breakfast some of us worked in the Science Dome. Clare demonstrated for Rajvi how to test rocks for carbonates. Daniel carried out another PCR and gel visualisation for ID of the DNA that was extracted from a second soil sample, which Rajvi assisted with. This found Archaea and Fungi in sample. More media dilutions of the brought 5 different media were prepared and inoculated with the soil and “yellow goo” collected yesterday. A non-targeted PCR was run to identify if there was any DNA present in the “yellow goo” and soil put into media in the morning. Meanwhile I tried to draw a sample of honeycomb weathering. Bharti planned her base geology map and wrote the introduction to her paper on the geology of the area. She is currently working on methodology. Aditya continued analysis of a new image of the Crab Nebula. Mehnaz gardened and worked on data cleaning and her codes. This afternoon Clare, Bharti and I left in Curiosity and Opportunity for EVA #11 to Candor Chasma. Our EVA was three and a half hours long and we left the rovers parked and walked about a kilometre to the canyon. At the chasma we climbed down, noticing the bedding plane, the contact between Salt Wash and the Sommerville formations and the geomorphology of the canyon. Then we traced the pathway of the dry riverbed from the ridge. We were home in time for a late afternoon cup of tea. Before dinner Aditya and Mehnaz did drawings in the science lab, exploring ideas about their research. This evening we are still busy, extracting DNA from rocks, all working on our research and reports (as this is our last week here) and drawing the dark sky. Pizza for dinner.

Look Ahead Plan: One EVA requested tomorrow to Candor Chasma and Compass Rock.

Anomalies in work: None

Weather: Cold but sunny all day 39F average with clouds early on and clear skies all afternoon.

Crew Physical Status: Nominal

EVA: EVA #11 to Candor Chasma.

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, Astronomy Report, EVA request.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site